# Neural Speech and Audio Coding

### v1

@@ 1 Introduction
Recently, data-driven methods for coding have gained intensive attention. In these neural speech and audio coding (NSAC) systems, a simple approach is to postulate an autoencoding task that reconstructs the input signal as exactly as possible in its output. In the meantime, to achieve the signal compression goal in the autoencoding pipeline, it is typical to decompose the system into two parts: an encoder module that produces a latent representation, followed by a decoder that recovers raw signal from the intermediate variable (Figure 2). Hence, it is critical to impose additional constraints on the encoder output so that it works as the code for transmission. For example, the code vector is supposed to be low in dimensionality and robust to quantization for the maximum coding gain. Such autoencoders can be trained in an end-to-end fashion with an adequately defined reconstruction loss as well as an additional entropy control mechanism. Kankanahali proposed a neural speech coding model [10] that consists of fully convolutional encoder and decoder modules, whose bottleneck feature map is discretized via the soft-to-hard quantization method [11]. Although this seminal end-to-end setup was successful to some degree, it fell short of competing with the legacy speech codec, e.g., AMR-WB [5]. More recently, such end-to-end autoencoder-based codecs have achieved higher coding gain, e.g., Soundstream [12], EnCodec [13], and Descript Audio Codec (DAC) [14], where residual vector quantization and adversarial loss functions play big roles. Another line of seminal work chose a powerful decoder architecture. A highly effective generative model, WaveNet [15], was employed as the decoder, which is trained to re-synthesize the speech signal by being indirectly conditioned on the very low-bitrate deterministic codes defined in the Codec 2 system [16] as shown in [17] or codes learned from VQ-VAE as in [18]. While these systems achieved phenomenal perceptual performance, they came at the cost of WaveNet’s high computational complexity.
@@ 2 Potentials and Limitations of NSAC
On the other hand, some neural speech codecs have shown sensible improvement in compression ratio while keeping the same level of speech quality with traditional codecs, i.e., from “Good” [17, 20] to “Excellent” [18] in MUSHRA [22] at the cost of higher complexity. Neural speech codecs’ very low bitrates can also improve their robustness to packet loss because the low bitrates allow for redundantly many packets for communication [23]. Another potential convenience in developing neural codecs is that it is relatively more straightforward to develop a universal codec that works for both speech and non-speech audio signals by training the model with various types of audio signals in a data-driven fashion. However, so far, the literature has shown that dedicated neural speech codecs can achieve better speech reconstruction than attempting to handle general audio signals. Indeed, general-purpose high-fidelity audio coding still remains a challenge for neural audio codecs despite substantial potentials shown in some neural audio codecs in low bitrates ($<$ 12 kbps) [12, 13, 14]. Note that the subjective listening test results reported in various NSAC papers are not directly comparable as they were done with different combinations of coding systems, adding more subjectivity to the individual results. Instead, in this paper we discuss their general trend. In addition, we opt to use the term NSAC to encompass both neural speech and audio codecs, because they can be potentially extended to cover both types of signals in the future.
@@ 4.2 LPCNet
A more sophisticated approach to combining the model-based and data-driven methods is LPCNet. It was originally proposed as a general-purpose vocoder, and then soon developed into a codec by being able to synthesize speech from a very low-bitrate (1.6 kbps) cepstrum-based code [20]. The main idea of LPCNet-based vocoding on the receiver side is, first, to estimate the LPC coefficients $a_{\tau}$ from the quantized cepstrum code $\bm{h}$. Secondly, the transmitter does not send the residual signal $\bm{e}$, although it is essential for LPC synthesis. Instead, LPCNet employs an RNN-based data-driven module that predicts the missing residual signal from the other available information within the decoder, such as the past samples generated so far $[\hat{x}_{t-p},\ldots,\hat{x}_{t-1}]$, LPC prediction $u_{t}$, and an auxiliary feature vector induced from the code $\bm{f}$ (Figure 4 for an illustration):
@@ 5 Learning to Predict Speech and Audio Signals in the Feature Space
$\bar{\bm{h}}_{t}\leftarrow\mathcal{W}^{\oplus}(\hat{\bm{h}}_{t},\bm{e}_{t}).$ (15)
@@ 5.1 Feature Prediction for the LPCNet Codec
$\bm{e}_{t}\leftarrow\bm{h}_{t}-\hat{\bm{h}}_{t},\quad\bar{\bm{h}}_{t}\leftarrow\hat{\bm{h}}_{t}+\bm{e}_{t}.$ (17)
@@ 6 Psychoacoustic Models for Perceptual Loss Functions
In MPEG audio coding, for example, MP3 or advanced audio coding (AAC), the bit allocation algorithm uses the noise-to-mask ratio (NMR). Given a masker tone’s logarithmic power spectrum density (PSD) in the $b$-th critical band $S_{b}$, the psychoacoustic model determines a masking threshold value $M_{f}$ in all subbands $f$ affected by the masker, which defines the signal-to-mask ratio (SMR), i.e., $\text{SMR}=S_{b}-M_{f}$. Meanwhile, if there are $m$ bits allocated for that subband $f$, which creates a noise level of $N_{f}$, the signal-to-noise ratio (SNR) is defined by $\text{SMR}(m)=S_{b}-N_{f}$. Eventually, NMR is defined as the difference (in decibels) between SNR and SMR,

### v2

@@ 1 Introduction
Recently, data-driven methods for coding have gained intensive attention. In these neural speech and audio coding (NSAC) systems, a simple approach is to postulate an autoencoding task that reconstructs the input signal as exactly as possible in its output. In the meantime, to achieve the signal compression goal in the autoencoding pipeline, it is typical to decompose the system into two parts: an encoder module that produces a latent representation, followed by a decoder that recovers raw signal from the intermediate variable (Figure 2). Hence, it is critical to impose additional constraints on the encoder output so that it works as the code for transmission. For example, the code vector is supposed to be low in dimensionality and robust to quantization for the maximum coding gain. Such autoencoders can be trained in an end-to-end fashion with an adequately defined reconstruction loss as well as an additional entropy control mechanism. Kankanahali proposed a neural speech coding model [10] that consists of fully convolutional encoder and decoder modules, whose bottleneck feature map is discretized via the soft-to-hard quantization method [11]. Although this seminal end-to-end setup was successful to some degree, it fell short of competing with the legacy speech codec, e.g., adaptive multi-rate wideband (AMR-WB) [5]. More recently, such end-to-end autoencoder-based codecs have achieved higher coding gain, e.g., Soundstream [12], EnCodec [13], and Descript Audio Codec (DAC) [14], where residual vector quantization and adversarial loss functions play big roles. Another line of seminal work chose a powerful decoder architecture. A highly effective generative model, WaveNet [15], was employed as the decoder, which is trained to re-synthesize the speech signal by being indirectly conditioned on the very low-bitrate deterministic codes defined in the Codec 2 system [16] as shown in [17] or codes learned from vector-quantized variational autoencoder (VQ-VAE) as in [18]. While these systems achieved phenomenal perceptual performance, they came at the cost of WaveNet’s high computational complexity.
@@ 2 Potentials and Limitations of NSAC
On the other hand, some neural speech codecs have shown sensible improvement in compression ratio while keeping the same level of speech quality with traditional codecs, i.e., from “Good” [17, 20] to “Excellent” [18] in multiple stimuli with hidden reference and anchor (MUSHRA) [22] at the cost of higher complexity. Neural speech codecs’ very low bitrates can also improve their robustness to packet loss because the low bitrates allow for redundantly many packets for communication [23]. Another potential convenience in developing neural codecs is that it is relatively more straightforward to develop a universal codec that works for both speech and non-speech audio signals by training the model with various types of audio signals in a data-driven fashion. However, so far, the literature has shown that dedicated neural speech codecs can achieve better speech reconstruction than attempting to handle general audio signals. Indeed, general-purpose high-fidelity audio coding still remains a challenge for neural audio codecs despite substantial potentials shown in some neural audio codecs in low bitrates ($<$ 12 kbps) [12, 13, 14]. Note that the subjective listening test results reported in various NSAC papers are not directly comparable as they were done with different combinations of coding systems, adding more subjectivity to the individual results. Instead, in this paper we discuss their general trend. In addition, we opt to use the term NSAC to encompass both neural speech and audio codecs, because they can be potentially extended to cover both types of signals in the future.
@@ 4.2 LPCNet
A more sophisticated approach to combining the model-based and data-driven methods is LPCNet. It was originally proposed as a general-purpose vocoder, and then soon developed into a codec by being able to synthesize speech from a very low-bitrate (1.6 kbps) cepstrum-based code [20]. The main idea of LPCNet-based vocoding on the receiver side is, first, to estimate the LPC coefficients $a_{\tau}$ from the quantized cepstrum code $\bm{h}$. Secondly, the transmitter does not send the residual signal $\bm{e}$, although it is essential for LPC synthesis. Instead, LPCNet employs an RNN-based data-driven module that predicts the missing residual signal from the other available information within the decoder, such as the past samples generated so far $[\hat{x}_{t-p},\ldots,\hat{x}_{t-1}]$, LPC prediction $u_{t}$, and an auxiliary feature vector an auxiliary feature vector $\bm{f}$ induced from the code $\bm{h}$ (Figure 4 for an illustration):
@@ 5 Learning to Predict Speech and Audio Signals in the Feature Space
$\bar{\bm{h}}_{t}\leftarrow\mathcal{W}^{\oplus}(\hat{\bm{h}}_{t},\bar{\bm{e}}_{t}).$ (15)
@@ 5.1 Feature Prediction for the LPCNet Codec
$\bm{e}_{t}\leftarrow\bm{h}_{t}-\hat{\bm{h}}_{t},\quad\bar{\bm{h}}_{t}\leftarrow\hat{\bm{h}}_{t}+\bar{\bm{e}}_{t}.$ (17)
@@ 6 Psychoacoustic Models for Perceptual Loss Functions
In MPEG audio coding, for example, MP3 or advanced audio coding (AAC), the bit allocation algorithm uses the noise-to-mask ratio (NMR). Given a masker tone’s logarithmic power spectrum density (PSD) in the $b$-th critical band $S_{b}$, the psychoacoustic model determines a masking threshold value $M_{f}$ in all subbands $f$ affected by the masker, which defines the signal-to-mask ratio (SMR), i.e., $\text{SMR}=S_{b}-M_{f}$. Meanwhile, if there are $m$ bits allocated for that subband $f$, which creates a noise level of $N_{f}$, the signal-to-noise ratio (SNR) is defined by $\text{SNR}(m)=S_{b}-N_{f}$. Eventually, NMR is defined as the difference (in decibels) between SNR and SMR,